# How to Find Fantastic Papers: Self-Rankings as a Powerful Predictor of Scientific Impact Beyond Peer Review

### v1

@@ 7 Discussion
The robustness of these findings is supported by analyses across multiple impact metrics. The predictive power of self-rankings held true not only for citations tracked by Semantic Scholar but also for Google Scholar citations and for a community engagement metric, GitHub stars. While these results are compelling, we acknowledge the need for validation over a longer time horizon and through experiments at more conferences to better assess the long-term impact of research work. To this end, we have already conducted similar experiments at ICML 2024 and 2025, as well as NeurIPS 2025, which we expect will validate and extend these findings. Further challenges include ensuring truthful reporting of self-rankings and determining how self-rankings can be formally incorporated into the review process. Notably, the comparative design of self-rankings renders it impossible for authors to simply inflate the quality of all their submissions. Indeed, an emerging body of literature is developing mechanisms in the game-theoretic framework that leverage rankings for more accurate review scores (Su, 2021; Yan et al., 2025).
These results carry significant implications for the future of scholarly evaluation, particularly in fast-moving fields like AI. Such fields demand review systems capable of separating methodological correctness from likely scientific consequence. Current peer review in AI research tends to reward correctness and quantifiable improvements over broader impact. The increasing use of language models in the review process may further bias evaluations toward surface-level characteristics rather than profound scientific contributions. In contrast, our study demonstrates that self-rankings appear better aligned with anticipated field-shaping value, perhaps because authors possess a deeper, more holistic understanding of their work’s conceptual novelty and long-term potential. This information is particularly valuable in the selection of best paper awards at AI conferences, which are intended to promote high-impact research. While more research is needed to determine precisely how to integrate authors’ self-rankings into review processes, our work provides compelling evidence that this information from authors is a valuable, low-cost, and reliable signal for identifying the research that will shape the future.
@@ Acknowledgments
Appendix A Details on the experiment
A.1 Survey questions
Figure S.1 presents screenshots of the surveys. Authors with multiple submissions were asked to rank their papers based on their perceived quality by dragging them up or down. All authors were asked to respond to the questions shown in Figure S.1. A comprehensive analysis of survey questions other than the self-ranking is presented in (Su et al., 2025).
@@ Appendix B Additional analyses
B.1 Additional details for Sections 3-5
@@
B.2 Additional details for Section 6
@@
We also note that, in the upper right panel of Figure S.4, to quantify the posting time, we converted each arXiv posting datetime into the number of seconds since January 1, 1970 at 00:00:00 UTC, commonly known as the UNIX timestamp. A Welch’s $t$-test comparing the UNIX timestamp in two groups yields a $p$-value of $0.72$, suggesting that arXiv posting time does not significantly differ between groups and is unlikely to confound our analysis. Moreover, higher proportion of the low-ranked papers posted either very early or very late compared to the high-ranked papers. This could be attributed to differences in submission quality. Low-ranked papers are more likely to have been rejected from prior venues and later recycled for submission to ICML 2023 (see Table 1 in Su et al. (2025)). Similarly, papers posted very late on arXiv often correspond to submissions that were rejected from ICML 2023 and later resubmitted to subsequent AI conferences. Finally, we make a observation related to the adversarial aspects of the double-blind review process. Notably, over 25% of submissions were posted on arXiv before the paper submission deadline, and over 50% were posted prior to the review release to authors. Such early postings may introduce bias into the review process and pose challenges to maintaining authors’ anonymity.
@@
B.3 Additional details for Section 7

### v2

@@ 7 Discussion
Citation counts are usually viewed as a strong indicator of a paper’s impact in AI. At AI conferences, papers typically introduce fresh ideas—either novel insights into existing models or entirely new methods—and there isn’t a single “ground truth” that immediately proves which approach is best; instead, the community tests, debates, and builds on them. In that context, a paper’s citation count serves as a practical signal of scientific impact: when many researchers cite a paper, it means the idea has been noticed, scrutinized, and found useful enough to inform further work, making it a reasonable proxy for scientific influence. Beyond academia, AI papers can also shape products and markets. Popular, highly cited ideas tend to attract developer mindshare, funding, and tooling—so the “hotter” the idea, the more people try it and the greater its perceived business value. That’s why heavily cited architectures (e.g., the Transformer (vaswani2017attention)) have become the backbone of many commercial systems.
Beyond citation counts, the robustness of our findings is supported by analyses across multiple impact metrics. The predictive power of self-rankings held true not only for citations tracked by Semantic Scholar but also for Google Scholar citations and for a community engagement metric, GitHub stars. While these results are compelling, we acknowledge the need for validation over a longer time horizon and through experiments at more conferences to better assess the long-term impact of research work. To this end, we have already conducted similar experiments at ICML 2024 and 2025, as well as NeurIPS 2025, which we expect will validate and extend these findings. Further challenges include ensuring truthful reporting of self-rankings and determining how self-rankings can be formally incorporated into the review process. Notably, the comparative design of self-rankings renders it impossible for authors to simply inflate the quality of all their submissions. Indeed, an emerging body of literature is developing mechanisms in the game-theoretic framework that leverage rankings for more accurate review scores (su2021you; yan2023isotonic).
These results carry significant implications for the future of scholarly evaluation, particularly in fast-moving fields like AI. Such fields demand review systems capable of separating methodological correctness from likely scientific consequences. Current peer review in AI research tends to reward correctness and quantifiable improvements over broader impact. The increasing use of language models in the review process may further bias evaluations toward surface-level characteristics rather than profound scientific contributions. In contrast, our study demonstrates that self-rankings appear better aligned with anticipated field-shaping value, perhaps because authors possess a deeper, more holistic understanding of their work’s conceptual novelty and long-term potential. This information is particularly valuable in the selection of best paper awards at AI conferences, which are intended to promote high-impact research. Our work provides compelling evidence that authors’ self-rankings are a low-cost, reliable signal that can be incorporated into future peer review systems. Indeed, based on the findings from the ICML 2023 experiment, ICML 2026 will, for the first time, incorporate author self-rankings into its formal decision-making process (see https://icml.cc/Conferences/2026/CallForPapers).
Methods
Methods are available in the Section A.
@@ Acknowledgments
Appendix A Methods
A.1 Survey Question
Figure S.1 presents full screenshots of the surveys. Authors with multiple submissions were asked to rank their papers based on their perceived quality by dragging them up or down. All authors were asked to respond to the questions shown in Figure S.1. A comprehensive analysis of survey questions other than the self-ranking is presented in (su2025icml).
@@ Appendix B Additional analyses
B.1 Additional details of Section 3-5
@@
B.2 Additional details of Section 6
@@
We also note that, in the upper right panel of Figure S.4, to quantify the posting time, we converted each arXiv posting datetime into the number of seconds since January 1, 1970 at 00:00:00 UTC, commonly known as the UNIX timestamp. A Welch’s $t$-test comparing the UNIX timestamp in two groups yields a $p$-value of $0.72$, suggesting that arXiv posting time does not significantly differ between groups and is unlikely to confound our analysis. Moreover, higher proportion of the low-ranked papers posted either very early or very late compared to the high-ranked papers. This could be attributed to differences in submission quality. Low-ranked papers are more likely to have been rejected from prior venues and later recycled for submission to ICML 2023 (see Table 1 in su2025icml). Similarly, papers posted very late on arXiv often correspond to submissions that were rejected from ICML 2023 and later resubmitted to subsequent AI conferences. Finally, we make a observation related to the adversarial aspects of the double-blind review process. Notably, over 25% of submissions were posted on arXiv before the paper submission deadline, and over 50% were posted prior to the review release to authors. Such early postings may introduce bias into the review process and pose challenges to maintaining authors’ anonymity.
@@
B.3 Additional details of Section 7